OpenStudy (anonymous):

what is the answer to ... -4 - (-3) + 5

OpenStudy (anonymous):

-4+8+5 -4+13 9

OpenStudy (anonymous):

=-4+3+5 =-4+8 =4

OpenStudy (anonymous):

4

OpenStudy (anonymous):

I got -2

OpenStudy (anonymous):

im a life saver now!! lol.. thank you!! :) / cookie how?

OpenStudy (anonymous):

cookie, the double negative sign becomes addition, check your work there

OpenStudy (anonymous):

- 4 - (-3) + 5 = - 4 + 3 = - 7 + 5 <<<< negative signs become subtraction = -2 that's how I did it

OpenStudy (anonymous):

the second step, -4 + 3 would become -1, and then -1 plus 5 is 4

OpenStudy (anonymous):

* I mean opposite signs become subtraction

OpenStudy (anonymous):

-4+3= -1 ..

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

yes^

OpenStudy (anonymous):

I forgot all about that Thank-you guys ! :D

OpenStudy (anonymous):

no problem, good luck